I need you to help with a graphical method. I do not understand it or algebra to beging with. I am stuck taking this class. Please help me solve some equations in my text book. Please explain how I can come to the conclusion of the answer so I can better understand it.
These are the current problems I am working on.
1. 3x plus 2y=0 I need to show the graphical method and how I got to the solution. But I am having trouble understanding the method.
2. the next question that I do not understand is about the system of equations by the method of substitution. 3x minus 7y=10
negitive 2x+y=negitive 14
I do not understand what I am suppose to do.
3. Another problem that I don't get is to solve the system of the linear equation by the method of elimination. 3x+2y=11
x minus 3y=negitive 11
4. I need to make a graph of the solution of the system of linear inequalities. 1/2x+y>4
x <6
y<3
5. I need to solve the system of linear equations using the method of Gaussian elimination.
-x+y+2z=1
2x+3y+z=-2
5x+4y+2z=4
I thank you very much for your help with these. I am complete lost, and was hoping this would give me some insite on how to do these problems. Please explain to me in detail of how to get to the answer so I know for further reference.























































































































Answer by mananth(16949)

1. 3x + 2y=0
You have to generate values for y from x
if x=2, y=-3 (2,-3)
if x= -2 , y = y=2 (-2,3)
if x=4, y = -4, ( (4,-6)
you have generated co-ordinates of three points.
plot the points on a graph and draw the straight line joining them


You have to find the value of a and y.
I am not too sure whether i could explain
3 x -7 y = 10 .............1
Total value
-2 x + 1 y = -14 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 7
3 x -7 y = 10
-14 x + 7 y = -98
Add the two equations
-11 x = -88
/ -11
x = 8
plug value of x in (1)
3 x -7 y = 10
24 -7 y = 10
-7 y = 10 -24
-7 y = -14
y = 2
x= 8
y= 2
m.ananth@hotmail.ca